How much water must be evaporated from 8 gal of an 8% salt solution in order to obtain a 12% salt solution?
Answer by Alan3354(21605)

How much water must be evaporated from 8 gal of an 8% salt solution in order to obtain a 12% salt solution?
--------------
At 8%, it's 0.64 gallons of salt and 7.36 gallons of water.
For the 0.64 gallons of salt to be 12%, the total must be 0.64/0.12
0.64/0.12 = 16/3 gallons total to remain.
8 - 16/3 must evaporate
= 8/3 gallons to evaporate